Solve this by graphing: x^2 +2x - 8 = 0
I had tried to graph this for several hours, but I still can't figure it out.

Answer by ewatrrr(24785)

 
Hi,
y = x^2 +2x - 8
y = x^2 +2x +1 - 1 - 8
y = (x+1)^2 - 9 V(-1,-9) and line of symmetry is x = -1
and when y = 0 ⇒ (x+1)^2 = 9, x = -1 ± 3, Pt(2,0) & (-4,0) are the x-intercepts